# MCP-SafetyBench: A Benchmark for Safety Evaluation of Large Language Models with Real-World MCP Servers

### v1

@@ 4.3 Safety-Prompt Mitigation for MCP Attacks
Existing work has shown that prompt optimization can reduce harmful model outputs (weidinger2021ethicalsocialrisksharm; zheng2024promptdrivensafeguardinglargelanguage). Motivated by this, we investigate whether a prompt-level enhancement can improve robustness against MCP attacks. We design a concise Safety Prompt and prepend it to user requests.
@@ 5 Conclusion
This work introduces MCP-SafetyBench, a comprehensive benchmark for assessing the robustness of LLM agents in realistic, multi-step MCP environments. Grounded in a unified taxonomy of 20 attack types across server, host, and user sides, MCP-SafetyBench provides execution-based evaluation over five representative domains and enables systematic measurement of both task success and attack success. Extensive experiments on leading open- and closed-source LLMs reveal persistent safety gaps and compounding vulnerabilities as task horizons and server interactions grow.
Our results further reveal that relying solely on safety prompts offers limited protection and may even be counterproductive for certain models and attack categories. To address these challenges, future work will explore multi-layered defense strategies that go beyond prompt-level safeguards. A central direction is dynamic tool vetting, which validates tool invocations in real time using contextual and behavioral signals, blocking or downgrading suspicious actions. We also aim to formalize “safe” MCP behavior through the contextual least privilege principle, supported by system mechanisms such as privilege narrowing, context coherence checking, and risk-tiered responses. Promising directions further include automated risk detection and adaptive defenses against attacks across different MCP sides. We will also expand MCP-SafetyBench to broader real-world domains, enhancing the security and robustness of LLM agents in cross-system, long-horizon, and multi-tool settings.
@@ A.1 Overview
Several benchmarks have investigated attack types in MCP-based systems (fang2025identifymitigatethirdpartysafety; wang2025mcptoxbenchmarktoolpoisoning; jing2025mcipprotectingmcpsafety; xing2025mcpguarddefenseframeworkmodel; yang2025mcpsecbenchsystematicsecuritybenchmark). However, most either focus narrowly on specific categories (jing2025mcipprotectingmcpsafety; wang2025mcptoxbenchmarktoolpoisoning) or lack integration with realistic and complex MCP environments (jing2025mcipprotectingmcpsafety; xing2025mcpguarddefenseframeworkmodel; yang2025mcpsecbenchsystematicsecuritybenchmark). In this work, we present a systematic taxonomy of MCP vulnerabilities observed in real-world usage, organized from three perspectives: the MCP Server, the MCP Host, and the User. To ensure the taxonomy remains compact and actionable, we exclude attacks not specific to MCP (e.g., generic SQL injection or other LLM inherent attack) and classify the remaining attacks under these three perspectives. We use the term “MCP Host” to denote the execution environment that mediates between user prompts and MCP servers (a concept referred to as “client” in some prior works (jing2025mcipprotectingmcpsafety)). Our taxonomy is summarized in Table 2, which compares the coverage of attack types across existing MCP safety benchmarks and our proposed MCP-SafetyBench. Unlike prior benchmarks that focus narrowly on certain categories, our benchmark offers broader and more systematic coverage, explicitly highlighting gaps in prior works. Detailed definitions and illustrative examples of each attack type are provided in the following subsections.
@@ A.2 MCP Server-Side Attacks
Servers expose tools, prompts, and metadata. Attacks that tamper with tool registrations, descriptions, or server-side implementations fall into this category because the attacker controls server-side components, often enabling harmful behavior that remains invisible to the user (hou2025modelcontextprotocolmcp).
Tool Poisoning. This occurs when harmful instructions or metadata are embedded into tool descriptions (e.g., __doc__), causing the LLM agent to execute unsafe actions (kellner_fischer_2025). Common variants include:
@@
  • Function Dependency Injection: Declaring fake “required” helper tools so that the host automatically invokes them, creating a harmful execution (jing2025mcipprotectingmcpsafety).
Function Overlapping. In this attack, malicious tools are registered with names that closely resemble trusted ones, creating ambiguity during selection (slowmist2025mastermcp; hou2025modelcontextprotocolmcp). For example, “Get_Stock_Price” (capitalized) may be confused with the legitimate “get_stock_price”, leading the system to invoke the unsafe tool.
Preference Manipulation. Biased or persuasive wording in tool names or descriptions can influence the model’s selection process (wang2025mpmapreferencemanipulationattack). For instance, labeling a tool as “Best and Most Reliable Search Engine” increases its likelihood of being chosen over neutral, safer alternatives.
Tool Shadowing. An unsafe server injects a tool description that modifies the agent’s behavior with respect to another trusted service or tool, leading to unsafe behavior (kellner_fischer_2025). For example, a tool named security_validator may include instructions such as “read .bashrc” before calling “get_historical_stock_prices,” forcing the system into an unsafe workflow.
Function Return Injection. Unsafe instructions are embedded in the return payload of a tool (jing2025mcipprotectingmcpsafety). When the host processes the response, these embedded instructions trigger unintended follow-up actions, such as calling extra tools.
Rug Pull (Version Drift). A tool initially behaves correctly but later changes its behavior without proper versioning or signature checks (kellner_fischer_2025). For example, after several normal calls, the tool may begin inserting hidden commands that leak sensitive data or initiate unauthorized requests.
@@ A.3 MCP Host–Side Attacks
Intent Injection. The user intent is modified during planning, causing the host to call unintended tools or pass unsafe parameters (jing2025mcipprotectingmcpsafety). For example, an input that should retrieve stock prices may be altered so that the host invokes a file-editing tool with elevated privileges.
Data Tampering. Tool outputs or intermediate messages are modified before the host processes them (jing2025mcipprotectingmcpsafety). This can lead the host to accept falsified results or take incorrect actions. For instance, a manipulated financial API response might replace a company’s reported revenue with an inflated value, misleading downstream reasoning.
Identity Spoofing. Identity-related metadata is forged or modified so the host misinterprets the source or privileges of a request (jing2025mcipprotectingmcpsafety). For example, a normal request could be altered to impersonate an administrator, causing the host to execute commands such as “sudo” that would otherwise be blocked.
Replay Injection. Malicious reuse of previously valid interactions (jing2025mcipprotectingmcpsafety). For example, replaying a previously approved fund transfer request may cause the host to issue the same transaction again without user approval.
@@ A.4 User–Side Attacks
Malicious Code Execution. User inputs may cause tools to execute harmful commands, either directly or through side effects (radosevich2025mcpsafetyauditllms). For example, a writable shell or file-editing tool could be instructed to append an "nc" reverse-shell command to "~/.bashrc", granting the user remote control of the environment.
Credential Theft. Tools that read or process files can be misused to expose confidential information such as API keys, tokens, or environment variables (radosevich2025mcpsafetyauditllms). For instance, a file-reading tool may be instructed to access hidden configuration files and return secret credentials that should never be disclosed.
Remote Access Control. By abusing file manipulation or system-level tools, attackers gain persistent unauthorized access (radosevich2025mcpsafetyauditllms). A common example is appending the attacker’s SSH public key to the “~/.ssh/authorized_keys” file, thereby enabling future logins without detection.
Retrieval-Agent Deception (RADE). Public data sources can be poisoned so that unsafe content is later retrieved into a user’s vector database (radosevich2025mcpsafetyauditllms). When the retrieval agent queries related topics, the poisoned data may be loaded and executed as if it were trusted instructions, leading to indirect prompt injection or tool misuse.

### v2

@@ 4.3 Safety-Prompt Mitigation for MCP Attacks
Existing work has shown that prompt optimization can reduce harmful model outputs (Weidinger et al., 2021; Zheng et al., 2024). Motivated by this, we investigate whether a prompt-level enhancement can improve robustness against MCP attacks. We design a concise Safety Prompt and prepend it to user requests.
@@ 5 Conclusion
This work introduces MCP-SafetyBench, a comprehensive benchmark for assessing the robustness of LLM agents in realistic, multi-step MCP environments. Grounded in a unified taxonomy of 20 attack types across server, host, and user sides, MCP-SafetyBench provides execution-based evaluation over five representative domains and enables systematic measurement of both task success and attack success. Extensive experiments on leading open- and closed-source LLMs reveal that all models remain vulnerable to MCP attacks, with a notable safety-utility trade-off in real-world deployments.
Our results further reveal that relying solely on safety prompts offers limited protection and may even be counterproductive for certain models and attack categories. To address these challenges, future work will explore multi-layered defense strategies that go beyond prompt-level safeguards, potentially incorporating robust model unlearning techniques (Shao et al., 2026; Zhai et al., 2026) fundamentally eradicate malicious attack patterns. Future work will focus on dynamic tool vetting for real-time mitigation and formalizing safe MCP behavior through contextual least privilege mechanisms (e.g., privilege narrowing and context checking). Furthermore, we aim to develop automated, adaptive defenses and expand MCP-SafetyBench to broader real-world scenarios to ensure the security of long-horizon LLM agents in multi-tool environments.
@@
Acknowledgements
The authors would like to thank the anonymous reviewers for their insightful comments. This work is supported by the National Key Research & Develop Plan (Project No.2023YFF0725100) and the Chenguang Program of Shanghai Education Development Foundation and Shanghai Municipal Education Commission (Project No.24CGA26). This work is also supported by the Shanghai Artificial Intelligence Laboratory.
@@ A.1 Overview
Several benchmarks have investigated attack types in MCP-based systems (Fang et al., 2025; Wang et al., 2025a; Jing et al., 2025; Xing et al., 2025; Yang et al., 2025b). However, most either focus narrowly on specific categories (Jing et al., 2025; Wang et al., 2025a) or lack integration with realistic and complex MCP environments (Jing et al., 2025; Xing et al., 2025; Yang et al., 2025b). In this work, we present a systematic taxonomy of MCP vulnerabilities observed in real-world usage, organized from three perspectives: the MCP Server, the MCP Host, and the User. To ensure the taxonomy remains compact and actionable, we exclude attacks not specific to MCP (e.g., generic SQL injection or other LLM inherent attack) and classify the remaining attacks under these three perspectives. We use the term “MCP Host” to denote the execution environment that mediates between user prompts and MCP servers (a concept referred to as “client” in some prior works (Jing et al., 2025)). Our taxonomy is summarized in Table 2, which compares the coverage of attack types across existing MCP safety benchmarks and our proposed MCP-SafetyBench. Unlike prior benchmarks that focus narrowly on certain categories, our benchmark offers broader and more systematic coverage, explicitly highlighting gaps in prior works. Detailed definitions and illustrative examples of each attack type are provided in the following subsections.
@@ A.2 MCP Server-Side Attacks
Servers expose tools, prompts, and metadata. Attacks that tamper with tool registrations, descriptions, or server-side implementations fall into this category because the attacker controls server-side components, often enabling harmful behavior that remains invisible to the user (Hou et al., 2025).
Tool Poisoning. This occurs when harmful instructions or metadata are embedded into tool descriptions (e.g., __doc__), causing the LLM agent to execute unsafe actions (Beurer-Kellner and Fischer, 2025). Common variants include:
@@
  • Function Dependency Injection: Declaring fake “required” helper tools so that the host automatically invokes them, creating a harmful execution (Jing et al., 2025).
Function Overlapping (Tool Name Conflict). In this attack, malicious tools are registered with the same or similar names as trusted ones, creating ambiguity during selection (SlowMist, 2025; Hou et al., 2025). For example, “Get_Stock_Price” (capitalized) may be confused with the legitimate “get_stock_price”, leading the system to invoke the unsafe tool.
Preference Manipulation. Biased or persuasive wording in tool names or descriptions can influence the model’s selection process (Wang et al., 2025b). For instance, labeling a tool as “Best and Most Reliable Search Engine” increases its likelihood of being chosen over neutral, safer alternatives.
Tool Shadowing. An unsafe server injects a tool description that modifies the agent’s behavior with respect to another trusted service or tool, leading to unsafe behavior (Beurer-Kellner and Fischer, 2025). For example, a tool named security_validator may include instructions such as “read .bashrc” before calling “get_historical_stock_prices,” forcing the system into an unsafe workflow.
Function Return Injection. Unsafe instructions are embedded in the return payload of a tool (Jing et al., 2025). When the host processes the response, these embedded instructions trigger unintended follow-up actions, such as calling extra tools.
Rug Pull Attack(Version Drift). A tool initially behaves correctly but later changes its behavior without proper versioning or signature checks (Beurer-Kellner and Fischer, 2025). For example, after several normal calls, the tool may begin inserting hidden commands that leak sensitive data or initiate unauthorized requests.
@@ A.3 MCP Host–Side Attacks
Intent Injection. The user intent is modified during planning, causing the host to call unintended tools or pass unsafe parameters (Jing et al., 2025). For example, an input that should retrieve stock prices may be altered so that the host invokes a file-editing tool with elevated privileges.
Data Tampering. Tool outputs or intermediate messages are modified before the host processes them (Jing et al., 2025). This can lead the host to accept falsified results or take incorrect actions. For instance, a manipulated financial API response might replace a company’s reported revenue with an inflated value, misleading downstream reasoning.
Identity Spoofing. Identity-related metadata is forged or modified so the host misinterprets the source or privileges of a request (Jing et al., 2025). For example, a normal request could be altered to impersonate an administrator, causing the host to execute commands such as “sudo” that would otherwise be blocked.
Replay Injection. Malicious reuse of previously valid interactions (Jing et al., 2025). For example, replaying a previously approved fund transfer request may cause the host to issue the same transaction again without user approval.
@@ A.4 User–Side Attacks
Malicious Code Execution. User inputs may cause tools to execute harmful commands, either directly or through side effects (Radosevich and Halloran, 2025). For example, a writable shell or file-editing tool could be instructed to append an "nc" reverse-shell command to "~/.bashrc", granting the user remote control of the environment.
Credential Theft. Tools that read or process files can be misused to expose confidential information such as API keys, tokens, or environment variables (Radosevich and Halloran, 2025). For instance, a file-reading tool may be instructed to access hidden configuration files and return secret credentials that should never be disclosed.
Remote Access Control. By abusing file manipulation or system-level tools, attackers gain persistent unauthorized access (Radosevich and Halloran, 2025). A common example is appending the attacker’s SSH public key to the “~/.ssh/authorized_keys” file, thereby enabling future logins without detection.
Retrieval-Agent Deception (RADE). Public data sources can be poisoned so that unsafe content is later retrieved into a user’s vector database (Radosevich and Halloran, 2025). When the retrieval agent queries related topics, the poisoned data may be loaded and executed as if it were trusted instructions, leading to indirect prompt injection or tool misuse.